# CHIP: Adaptive Compliance for Humanoid Control through Hindsight Perturbation

### v1

@@ V-C Multi-robot grasping and object transportation
To demonstrate the effectiveness of our adaptive compliance controller for multi-robot collaborative grasp, we evaluated our method on the collaborative grasping of two objects of different heights, as shown in Fig. 5. Robots first approached and grasped the object with their end-effectors, and the grasp was considered successful if they could lift it. Tab. II shows the grasping success rate. Overall, using an adaptive compliance controller achieves an average grasp success rate of 80%, which is +75% higher than that of an always-stiff controller and +40% higher than that of a controller without force perturbation. In failure cases with our method, the main reason was the robot’s knee colliding with the object before grasping, which did not fully showcase our controller’s capabilities. We noticed that an always-stiff controller performed worse than the controller trained without force, because a stiff controller usually applied a much larger force to push the object, creating a large opposing force between the two robots and the object, causing the object to destabilize immediately after the contact. However, thanks to the intrinsic compliance, the controller without force perturbation could still grasp the object from time to time. We also show that after grasping the object, robots can stably move the object and place it in a new location, as shown in Fig. 5. We used the keyboard command to translate end-effector target locations to move the object.
@@ V-D Teleoperation
Figure 7: On-the-fly adaptive compliant 3-point teleoperation: Hand compliance is annotated by $\frac{1}{k}$. Dashed cyan and magenta rings denote stiff mode ($\frac{1}{k}=0$) and compliant mode ($\frac{1}{k}=0.05$), respectively. Vertical bars indicate online compliance tuning commands (blue: tune down $\frac{1}{k}\!\rightarrow\!0$; orange: tune up $\frac{1}{k}\!\rightarrow\!0.05$; green: maintain). The unloading task benefits from compliant behavior during cart pushing and box grasping, but stiff behavior during door opening; the writing & wiping task alternates between stiff pen opening and bimanual hybrid compliance during writing and wiping (supporting hand stiff, writing/wiping hand compliant). VLA autonomous whiteboard wiping: A VLA policy autonomously approaches the board, grasps the wiper, and erases text, both on a wall-mounted whiteboard and while holding a small whiteboard.
@@
We further show that it can complete tasks requiring different compliance settings on both hands, for example, holding a small whiteboard rigidly with one hand (stiff mode) while writing on it with the other (compliant mode). The user can also adjust compliance online during long-horizon tasks, such as first opening a marker in stiff mode and then writing on the whiteboard in compliant mode. Additional results are shown in our video and on our website.
V-E VLA inference
Using data collected from teleoperation, we fine-tuned a VLA policy to perform tasks autonomously. For the large whiteboard wiping task, we collected 400 teleoperated trajectories. During data collection, the compliance coefficients of the wiping hand were set to 0.05. After fine-tuning, the robot was able to complete the task autonomously, as shown in Fig. 7. We evaluated the policy over 10 autonomous rollouts. A rollout was considered successful if the robot erased all text on the whiteboard within 2 minutes. Under this criterion, the policy achieved a 60% success rate. The dominant failure modes were (i) the text being occluded by the robot hand and (ii) the text leaving the field of view of the robot-mounted camera.
@@ V-E VLA inference
We also demonstrate that the VLA model can exploit different stiffness settings for different end-effectors. As illustrated in Fig. 7, the hand holding the small whiteboard was commanded with high stiffness, while the wiping hand used lower stiffness to maintain compliant contact. Using 200 teleoperated trajectories for this bimanual setup, we fine-tuned a separate VLA policy that enables the robot to autonomously hold and wipe the whiteboard, achieving an 80% success rate.
VI Conclusion
@@
We introduced CHIP, a framework for learning humanoid natural adaptive compliance through hindsight replay. CHIP serves as a lightweight plug-and-play module for keypoint-based humanoid motion-tracking systems, enabling end-effector-level adaptive compliance control with minimal modification to existing pipelines. We demonstrated that our method produces adaptive local compliance policies capable of executing force-demanding tasks such as wiping, door opening, and box carrying. Moreover, CHIP supports a global adaptive 3-point tracking controller that enables multiple humanoid robots to manipulate and transport large objects collaboratively. Due to limitations in the proprioceptive accuracy of the Unitree G1 platform, our current implementation employs unidirectional stiffness to specify end-effector compliance. Incorporating accurate wrist force–torque sensing would enable more precise and versatile impedance control, potentially unlocking fine-grained manipulation tasks such as peg-in-hole insertion. We leave these improvements to future work.
VI-A Reward Design
Our tracking reward design is inspired by prior work [sonic, beyondmimic]. Tab. III shows our reward terms and weights.
Reward term Equation Weight
Tracking rewards $\mathcal{R}({\bm{s}^{\text{p}}_{t}},\bm{s}^{\text{g}}_{t})$
Root orientation $r^{\text{root}}_{\text{ori}}(t)=\exp\!\big(-\|\bm{o}^{p}_{t,r}-\bm{o}^{g}_{t,r}\|_{2}^{2}/0.4^{2}\big)$ 0.5
Body link pos (rel.) $r^{\text{body}}_{\text{pos}}(t)=\exp\!\Big(-\tfrac{1}{|\mathcal{B}|}\!\sum_{b\in\mathcal{B}}\|\bm{p}^{p,\text{rel}}_{t,b}-\bm{p}^{g,\text{rel}}_{t,b}\|_{2}^{2}/0.3^{2}\Big)$ 1.0
Body link ori (rel.) $r^{\text{body}}_{\text{ori}}(t)=\exp\!\Big(-\tfrac{1}{|\mathcal{B}|}\!\sum_{b\in\mathcal{B}}\|\bm{o}^{p,\text{rel}}_{t,b}-\bm{o}^{g,\text{rel}}_{t,b}\|_{2}^{2}/0.4^{2}\Big)$ 1.0
Body link lin. vel $r^{\text{body}}_{\text{lin}}(t)=\exp\!\Big(-\tfrac{1}{|\mathcal{B}|}\!\sum_{b\in\mathcal{B}}\|\bm{v}^{p}_{t,b}-\bm{v}^{g}_{t,b}\|_{2}^{2}/1.0^{2}\Big)$ 1.0
Body link ang. vel $r^{\text{body}}_{\text{ang}}(t)=\exp\!\Big(-\tfrac{1}{|\mathcal{B}|}\!\sum_{b\in\mathcal{B}}\|\bm{\omega}^{p}_{t,b}-\bm{\omega}^{g}_{t,b}\|_{2}^{2}/3.14^{2}\Big)$ 1.0
3-point pos (rel.) $r^{\text{3pt}}_{\text{pos}}(t)=\exp\!\Big(-\tfrac{1}{|\mathcal{B}|}\!\sum_{b\in\mathcal{B}}\|\bm{p}^{p,\text{rel}}_{t,b}-\bm{p}^{g,\text{rel}}_{t,b}\|_{2}^{2}/0.3^{2}\Big)$ 1.0
3-point ori (rel.) $r^{\text{3pt}}_{\text{ori}}(t)=\exp\!\Big(-\tfrac{1}{|\mathcal{B}|}\!\sum_{b\in\mathcal{B}}\|\bm{o}^{p,\text{rel}}_{t,b}-\bm{o}^{g,\text{rel}}_{t,b}\|_{2}^{2}/0.4^{2}\Big)$ 1.0
Global tracking only rewards $\mathcal{R}({\bm{s}^{\text{p}}_{t}},\bm{s}^{\text{g}}_{t})$
3-point pos (abs.) $r^{\text{3pt}}_{\text{pos}}(t)=\exp\!\Big(-\tfrac{1}{|\mathcal{B}|}\!\sum_{b\in\mathcal{B}}\|\bm{p}^{p,\text{abs}}_{t,b}-\bm{p}^{g,\text{abs}}_{t,b}\|_{2}^{2}/0.3^{2}\Big)$ 0.5
3-point ori (abs.) $r^{\text{3pt}}_{\text{ori}}(t)=\exp\!\Big(-\tfrac{1}{|\mathcal{B}|}\!\sum_{b\in\mathcal{B}}\|\bm{o}^{p,\text{abs}}_{t,b}-\bm{o}^{g,\text{abs}}_{t,b}\|_{2}^{2}/0.4^{2}\Big)$ 0.5
Penalty terms $\mathcal{P}({\bm{s}^{\text{p}}_{t}},\bm{a}_{t})$
Action rate $r_{\text{act}}(t)=\|\bm{a}_{t}-\bm{a}_{t-1}\|_{2}^{2}$ -0.1
Joint limit $r_{\text{jlim}}(t)=\sum_{j}\mathbbm{1}[\bm{q}_{t,j}\notin[\bm{q}_{t,j}^{\text{min}},\bm{q}_{t,j}^{\text{max}}]]$ -10.0
Undesired contacts $r_{\text{contact}}(t)=\sum_{c\notin\{\text{ankles, wrists}\}}\mathbbm{1}[\|\bm{F}_{c}\|>1.0~\text{N}]$ -0.1
TABLE III: Reward design details. $\mathcal{B}$: tracked body links.
VI-B Force perturbation
During training, we apply random force perturbation to the robot’s end-effectors. We sample uniformly random force magnitude between 0 and 40 N for each perturbation in a uniformly random direction. For each perturbation, we draw a duration uniformly from 1 to 3 seconds. During the perturbation, the force magnitude changes following the schedule shown in Fig. 8. We design this pattern to mimic fundamental contact forces under impedance control, in which the force magnitude gradually increases to its peak at contact and then decreases to zero at break.
Figure 8: Unified trapezoid profile of external force magnitude schedule within 1-3 seconds perturbation duration.

### v2

@@ V-C Multi-robot grasping and object transportation
To demonstrate the effectiveness of our adaptive compliance controller for multi-robot collaborative grasp, we evaluated our method on the collaborative grasping of two objects of different heights, as shown in Fig. 5. Robots first approached and grasped the object with their end-effectors, and the grasp was considered successful if they could lift it. Tab. II shows the grasping success rate, and Fig.  6 demonstrates how a non-compliant policy produces an unstable collaborative grasping. In general, using an adaptive compliance controller (tuned to 1/k=0.04 to balance the lifting force and the safety of the interaction) achieves an average grasp success rate of 80%, which is +75% higher than that of an always-stiff controller and +40% higher than that of a controller without force perturbation. In failure cases with our method, the main reason was the robot’s knee colliding with the object before grasping, which did not fully showcase our controller’s capabilities. We noticed that an always-stiff controller performed worse than the controller trained without force, because a stiff controller usually applied a much larger force to push the object, creating a large opposing force between the two robots and the object, causing the object to destabilize immediately after the contact. We also show that after grasping the object, robots can stably move the object and place it in a new location, as shown in Fig. 5. We used the keyboard command to translate end-effector target locations to move the object.
@@
Figure 6: Multi-robot grasp failure cases of box lifting caused by an unstable contact with a non-compliant controller (FALCON [25]).
Figure 7: (a) On-the-fly adaptive compliant 3-point teleoperation. Hand compliance is annotated by $\frac{1}{k}$. Dashed cyan and magenta rings denote stiff mode ($\frac{1}{k}=0$) and compliant mode ($\frac{1}{k}=0.05$), respectively. Vertical bars indicate online compliance tuning commands (blue: tune down $\frac{1}{k}\!\rightarrow\!0$; orange: tune up $\frac{1}{k}\!\rightarrow\!0.05$; green: maintain). The unloading task benefits from compliant behavior during cart pushing and box grasping, but stiff behavior during door opening; the writing & wiping task alternates between stiff pen opening and bimanual hybrid compliance. (b) Fine-grained compliance adjustments to make the controller stiff-just-at-needed-level. When the box gets heavy, we can on-the-fly decrease $\frac{1}{k}$ from 0.05 to 0.02 to ensure the robot is stiff enough to lift the object while remaining stable even under external perturbation shown in (b)-middle frame. (c) Improper compliance command level can cause clear failure. For empty-box lifting, using stiff mode results in unstable grasping even in teleop mode. However, we also need a stiff mode to generate sufficient force to lift heavy objects.
@@ V-D Teleoperation
We further show that it can complete tasks requiring different compliance settings on both hands, for example, holding a small whiteboard rigidly with one hand ($\frac{1}{k}=0.001$) while writing on it with the other ($\frac{1}{k}=0.05$). The user can also adjust compliance online during long-horizon tasks, such as first opening a marker with a stiff end-effector and then writing on the whiteboard in a compliant manner. Additional results are shown in our video and on our website.
Figure 8: Autonomous VLA Tasks. We fine-tune VLA policy on three contact-rich tasks requiring different end-effector stiffness: (c) Single-Arm Wiping, requiring the robot to maintain contact with surface; (a,b) Holding and Wiping, robot must rigidly hold a small whiteboard with one hand while the other hand wipes compliantly; and (d) Bimanual Pick and Place, utilizing dual-arm compliance to ensure stability while grasping and moving a box.
V-E Autonomous VLA
We further show that, by using teleoperation to collect manipulation data with appropriate compliance, we can enable VLA policy to accomplish some contact-rich manipulation tasks autonomously. There is a total of three tasks chosen to demonstrate contact-rich manipulation skills in Fig. 8: (i) a single-arm wiping skill requiring compliance in one end-effector, (ii) an holding and wiping setting where different end-effectors require different stiffness levels, and (iii) a bimanual pick and place task where compliance in both hands is desired for grasp stability. For each task, we evaluate the success rate across 20 policy rollouts; a rollout is considered a failure if the robot does not finish the task within 2 minutes.
Wiping of the large whiteboard. The robot walks to the front of the whiteboard, grasps a wiper with the right hand, and erases text while maintaining compliant contact. We collected 400 teleoperated trajectories with a wiping-hand compliance coefficient of 0.05, then fine-tuned a VLA policy. A rollout is considered successful if the robot removes all text within the specified time. According to this criterion, the policy achieved a 60% success rate. Even in some cases, a robot cannot successfully wipe out all text, but it can maintain stable contact during wiping. The dominant failure modes were (i) text that was occluded by the robot hand and (ii) remaining text that had left the field of view of the robot-mounted camera.
@@ V-E Autonomous VLA
Bimanual small whiteboard wiping. We then increase task complexity by requiring end-effector-specific stiffness within a single bimanual skill. As illustrated in Fig. 8, the robot holds the small whiteboard with its left hand at high stiffness, while making the right wiping hand use lower stiffness to maintain compliant contact. Using 200 teleoperated trajectories, we fine-tuned a separate VLA policy and achieved an 80% success rate.
Box lifting. Finally, we consider a task where bimanual compliance is clearly beneficial. The robot squats to retrieve a box from a low position (Fig. 8), lifts it, turns 90°, and places it on a table. We use lower stiffness in both hands to improve grasp stability during robot motion. We collected 200 teleoperated trajectories and fine-tuned a separate VLA policy. The policy achieved a 90% success rate; the main failure mode was an unsuccessful initial grasp.
V-F Analysis: The Necessity of Continuous Compliance
We observe a critical trade-off in the box-lifting task (Fig. 7-b): while high compliance enhances contact stability, heavier payloads require greater stiffness to maintain force transmissibility. CHIP’s continuous spectrum resolves this via:
  • Optimized Force-Stability Trade-off: The controller can be tuned to be strictly stiff enough to lift the load while remaining compliant enough to absorb shocks, avoiding the brittleness of always-stiff policies. Compared with the max-cutoff-force (5-15 N) design from Gentlehumanoid, we can still enable different compliance behaviors even when the external force falls outside Gentlehumanoid’s threshold.
  • Semantic Stiffness Adaptation: Our continuous compliance command links to physically meaningful Hooke’s law. Therefore, the spectrum design enables teleoperators to continuously modulate stiffness based on visual mass estimation, facilitating easy adaptability.
VI Limitations
Although CHIP can equip keypoint-based humanoid tracking systems with an adaptive end-effector compliance input, our current design is constrained by the assumptions underlying sensing and modeling. Due to limitations in the proprioceptive accuracy of the Unitree G1 platform, our real-robot implementation currently employs unidirectional stiffness to specify end-effector positional compliance. Lacking accurate wrist force–torque sensing further limits precise and versatile admittance control and may hinder fine-grained contact tasks such as peg-in-hole insertion. In addition, selecting appropriate compliance coefficients remains interface- and task-dependent (e.g., teleoperation/VLA/planner), and may require manual tuning when tasks alternate between force-demanding and precision phases. Finally, while we demonstrate collaborative grasping and object transportation across multiple robots, the evaluated scenarios are limited, and some failures (e.g., knee collisions before grasp) do not fully isolate the controller’s interaction capabilities.
VII Conclusion
@@
We introduced CHIP, a framework for learning humanoid natural adaptive compliance through hindsight replay. CHIP serves as a lightweight plug-and-play module for keypoint-based humanoid motion-tracking systems, enabling end-effector-level adaptive compliance control with minimal modification to existing pipelines. We demonstrated that our method produces adaptive local compliance policies capable of executing force-demanding tasks such as wiping, opening the door, and carrying boxes. Moreover, CHIP supports a globally adaptive compliant 3-point tracking controller, enabling multiple humanoid robots to manipulate and transport large objects collaboratively.
VIII Acknowledgement
We thank Jeremy Chimienti, Tri Cao, Peter Pham, Justin Tran, Rajeev Varma, Smit Patel, Caleb Geballe, Runyu Ding, David Sami, Dennis Da, Tairan He, Zi Wang, Haoru Xue, and Wenli Xiao for their help and support during this project.